Find the polynomial of degree three relevant to the following data using Lagrange's interpolation formula. Hence find f[4x0.(i+2)], where i= 2

X —> 1 3 5 6
y=f(x) —> -12 -14 -20 -20
1
Expert's answer
2021-08-27T02:07:49-0400

The degree of an individual term of a polynomial is the exponent of its variable; the exponents of the terms of this polynomial are, in order, 5, 4, 2, and 7. The degree of the polynomial is the highest degree of any of the terms; in this case, it is 7.

 The polynomial distribution:

f(x)=x^(2)+4,x>=0

6X5+5X4−3x2+x7

X —> 1 3 5

6 y=f(x) —> -12 -14 -20 -20